ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 228, 2018

el Abdalaoui, e. H., J. Kułaga-Przymus, M. Lemańczyk and T. de la Rue Möbius disjointness for models of an ergodic system and beyond

707

Abrams, G., F. Mantese and A. Tonolo Leavitt path algebras are Bézout

53

Accomazzo, N. A characterization of BMO in terms of endpoint bounds for commutators of singular integrals

787

Adolf, D., S. Cox and P. Welch Lower consistency bounds for mutual stationarity with divergent uncountable cofinalities

1

Arakelian, N. and H. Borges Bounds for the number of points on curves over finite fields

177

Berele, A. and A. Regev Growth of central polynomials of verbally prime algebras

201

Berestycki, N. and H. Jackson The Rohde–Schramm theorem via the Gaussian free field

973

Bik, A. and J. Draisma A note on ED degrees of group-stable subvarieties in polar representations

353

Borges, H. See N. Arakelian Checcoli, S. and J. Roques On some arithmetic properties of

Mahler functions

801

Cox, S. See D. Adolf Cuno, J. and E. Sava-Huss Random walks on Baumslag–Solitar

groups

627

Danilenko, A. I. Infinite measure preserving transformations with Radon MSJ

21

Debruyne, G. and J. Vindas Optimal Tauberian constant in Ingham’s theorem for Laplace transforms

557

Dolgopyat, D. and O. Sarig Asymptotic windings of horocycles  119

Doran, C. F., T. L. Kelly, A. Salerno, S. Sperber, J. Voight and U. Whitcher Zeta functions of alternate mirror Calabi–Yau families

665

Draisma, J. See A. Bik Felmer, P., D. dos Prazeres and E. Topp Interior regularity

results for zeroth order operators approaching the fractional Laplacian 835 Ginzburg, V. L., B. Z. Gürel and L. Macarini Multiplicity of

closed Reeb orbits on prequantization bundles

407

Gürel, B. Z. See V. L. Ginzburg Hu, H., E. Shi and Z. J. Wang Some ergodic and rigidity properties

of discrete Heisenberg group actions

933

Jackson, H. See N. Berestycki Kaliman, S., F. Kutzschebauch and T. T. Truong On subellip-

tic manifolds

229

Kang, M.-H., W.-C. W. Li and C. J. Wang Zeta and L-functions of finite quotients of apartments and buildings

79

Kelly, T. L. See C. F. Doran

Kot, P. Gradients of inner functions in strictly pseudoconvex domains  771

Krupiński, K., A. Pillay and T. Rzepecki Topological dynamics and the complexity of strong types

863

Kułaga-Przymus, J. See e. H. el Abdalaoui Kutzschebauch, F. See S. Kaliman Lemańczyk, M. See e. H. el Abdalaoui Li, W.-C. W. See M.-H. Kang Loiskekoski, L. and G. M. Ziegler Simple polytopes without small

separators, II: Thurston’s bound

293

Macarini, L. See V. L. Ginzburg

Malle, G. On a minimal counterexample to Brauer’s k(B)-conjecture  527

Mantese, F. See G. Abrams Paolini, G. and S. Shelah Polish topologies for graph products of

cyclic groups

305

Peluse, S. Three-term polynomial progressions in subsets of finite fields

379

Pillay, A. See K. Krupiński Praeger, C. E. and C. Schneider Group factorisations, uniform

automorphisms, and permutation groups of simple diagonal type

1001

dos Prazeres, D. See P. Felmer Regev, A. See A. Berele Roques, J. See S. Checcoli de la Rue, T. See e. H. el Abdalaoui Rzepecki, T. See K. Krupiński Salerno, A. See C. F. Doran

Sambale, B. Remarks on Harada’s Conjecture  513

Sarig, O. See D. Dolgopyat Sava-Huss, E. See J. Cuno Schneider, C. See C. E. Praeger

Sen, Y. Whitehead’s problem and reverse mathematics  455

Shelah, S. See G. Paolini Shi, E. See H. Hu Sperber, S. See C. F. Doran Stasinski, A. Commutators of trace zero matrices over principal ideal

rings

211

Thomas, S. Uniformly recurrent subgroups of simple locally finite groups

321

Tong-Viet, H. P. Real class sizes  753

Tonolo, A. See G. Abrams Topp, E. See P. Felmer

Tran, T. On the structure of large sum-free sets of integers  249

Truong, T. T. See S. Kaliman Vindas, J. See G. Debruyne Voight, J. See C. F. Doran Wang, C.-J. See M.-H. Kang Wang, Z. J. See H. Hu

Weaver, B. The local lifting problem for D_{4}  587

Welch, P. See D. Adolf Whitcher, U. See C. F. Doran Ziegler, G. M. See L. Loiskekoski